Please help me solve this problem:

Suppose your mom pays you to make a new dining room table for apartment. She tells you that the length of the table should be 3 feet more than twice its width. She also tells you that the area of the rectangular table top is 2w^2+12 square feet. What are the length and width of the table? Please explain!


thankyou so much!!!!!!!!

Answer by checkley77(12844)

L=2W+3
LW=2W^2+12
(2W+3)W=2W^2+12
2W^2+3W=2W^2+12
3W=12
W=12/3
W=4 FT. IS THE WIDTH.
L=2*4+3=8+3=11 FT. IS THE LENGTH.
PROOF:
4*11=2*4^2+12
44=2*16+12
44=32+12
44=44